OpenStudy (anonymous):

Check all that apply. If tan 0= 1/8, then: A. csc0 =17/15 B. sec0 =17/8 C. cot0 =8/15 D. cos0 =15/17

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

15/8 i mean

OpenStudy (anonymous):

oh lol ok

OpenStudy (anonymous):

tan = opp/adj opp = 15 adj = 8 15^2 + 8^2 = c^2 c= 17 Strange thing is A, B, and C are right D is wrong, should be 8/17

OpenStudy (anonymous):

then i mark a b and c

OpenStudy (anonymous):

Yep. I double checked :))

OpenStudy (anonymous):

that was right lol:D

OpenStudy (anonymous):

lol great xD first time see a question with 3 right answers though :P